five times the measure of the complemment of an angle is equal to 5/7 of the measure of the supplement of the angle. find the measure of the angle, its conplement, and its supplement.
thanks!

Answer by cleomenius(959)

I set this up as follows
5(90-x) = 5/7(180 - x)
450 - 5x = 128.6 - 5/7x
321.4 = 4 2/7x
x = app. 75 the angle we are working with.
So, now we check for the complement
90-75 = 15 the complement
15*5 = 75
Now we check for the supplement.
180 -75 = 105.
5/7 * 105 = 75.
This does check.
Cleomenius.